# An investigation of air pollution-induced temperature sensitivity and susceptibility to heat-related hospitalization in the Medicare population

### v1

@@ 3.2 Sensitivity analysis
To investigate the interaction between temperature and PM2.5 further, we developed an additional model using a tensor product to allow for a more flexible, nonlinear interaction between these exposures. This approach resulted in more linear estimates of the independent effects of temperature and PM2.5 compared to the main analysis (Figure S3 in the supplementary materials). The estimated RERI from this model was positive, but the 95% CI included zero and was smaller than the RERI estimated in the main analysis (Figures S4 and S5 and Table S2 in the supplementary materials). The wider credible intervals from the tensor model reflect increased uncertainty in the estimates when allowing for greater flexibility in the interaction structure.
@@ 4.1 Summary and key findings
To our knowledge, this is the first study to investigate the combined (and synergistic) effects of acute exposure to high temperatures and elevated PM2.5 on heat-related hospitalizations. Comparing median and 95th percentile exposure levels, we found a positive association between temperature and HRH (OR: 1.04; 95% CI: 1.03, 1.06). The association between PM2.5 and HRH was also positive, although the credible interval covered the odds ratio scale null value of one. Finally, we identified evidence of an additive interaction between temperature and PM2.5 (RERI: 0.03; 95% CI: 0.01, 0.06). Nevertheless, since temperature and PM2.5 are often correlated (see supplementary Figure S2), we have limited data to assess the effect of elevated PM2.5 at lower temperatures during the warm season.
Our findings offer the first empirical, epidemiological evidence that PM2.5 may increase risk of heat-related hospitalization. In contrast to prior work, which has largely framed temperature as a modifier of the health effects of air pollution (Li et al.,, 2017; Areal et al.,, 2022; Anenberg et al.,, 2020), our focus on explicit heat-related hospitalizations reveals that PM2.5 may also modify the health effects of heat.
@@ 4.2 Impact
Heat is the deadliest form of extreme weather in the United States (National Weather Service, 2024b, ), and rates of heat-related illness continue to climb (Vaidyanathan,, 2024; Osborne et al.,, 2023) even though these illnesses are predominantly preventable (Sorensen and Hess,, 2022; Gauer and Meyers,, 2019). Understanding how to protect vulnerable individuals, including older adults, from extreme heat events is a crucial component of climate change adaptation and an urgent public health issue (Vaidyanathan,, 2024).
@@
The National Weather Service issues localized heat alerts to initiate individual- and community-level preparations on unusually warm days (National Weather Service, 2024a, ). Our findings indicate that additional caution is necessary on hot days when PM2.5 levels are high, and it may be useful to consider PM2.5 when issuing heat alerts. However, the efficacy of heat alerts alone in minimizing illness and death is uncertain (Weinberger et al.,, 2018; Toloo et al.,, 2013). Previous research highlights that high-risk individuals must recognize their own vulnerabilities to heat, possess the knowledge and resources necessary to protect themselves, and take action to stay safe (Toloo et al.,, 2013). Interestingly, previous studies report that many older adults recognize age as a risk factor for heat-related illness but do not recognize high temperatures as a threat to their personal health (Widerynski et al.,, 2017). Clinicians could play an important role in helping these individuals understand their risk (Sorensen and Hess,, 2022).
@@
Our results also emphasize the importance of air conditioning, as open windows may allow more ambient particulate matter into homes (Lin et al.,, 2013) and further increase the risk of heat-related illness. These findings also contribute to the growing body of evidence that PM2.5 is linked to a range of negative health effects, underscoring the importance of reducing ambient PM2.5, particularly with a warming planet.
4.3 Biological plausibility
@@
This study supports mounting epidemiological evidence of synergism between temperature and PM2.5. Stafoggia et al., (2008) identified positive but generally non-statistically significant interactions between temperature and PM10, which includes both PM2.5 and larger particles, on mortality, and a meta-analysis by Li et al., (2017) also reported that extreme temperatures modify the effect of PM10 on non-accidental and cardiovascular mortality. Kioumourtzoglou et al., (2016) observed a stronger association between long-term PM2.5 exposure and mortality in warmer cities. A recent study also identified synergism between short-term exposures to temperature and PM2.5 for both respiratory and cardiac hospital admissions (Yitshak-Sade et al.,, 2018).
@@
Simultaneous exposure to heat and PM2.5 can exacerbate heat-related health outcomes through shared pathophysiological mechanisms, including autonomic dysfunction and heightened inflammatory responses. Both heat and PM2.5 cause autonomic dysfunction by increasing sympathetic drive and suppressing parasympathetic activity (Taylor-Clark,, 2020; Yamamoto et al.,, 2007). This imbalance leads to a 4%–7% reduction in heart rate variability and an increase in heart rate by 5–9 beats per minute (Shah et al.,, 2019; Florea and Cohn,, 2014; Jensen and Brabrand,, 2015). In addition, blood pressure increases due to autonomic vasoconstriction from PM2.5 exposure, with each 10 increment correlating with a 9 mm Hg rise in systolic pressure (Dvonch et al.,, 2009). Furthermore, increased sweating via sympathetic activation in response to heat elevates blood pressure by increasing plasma viscosity by 2% per C (Sakurai et al.,, 2004). Thus, the combined cardiovascular stress disrupts hemodynamic regulation, compromising the body’s thermoregulatory function and exacerbating heat stress.
@@
Heat and PM2.5 can also increase systemic inflammatory responses by forming reactive oxygen species within the cells. The resulting oxidative stress activates redox-sensitive transcription factors, driving the production and release of pro-inflammatory mediators such as tumor necrosis factor-alpha (TNF-), interleukin-6 (IL-6), and interleukin-1 (IL-1) (Sangkham et al.,, 2024; Ebi et al.,, 2021). A 10 increase in PM2.5 exposure corresponds to a fivefold rise in these cytokines, and a 2C elevation in ambient temperature can boost their expression by up to fourfold (Hu et al.,, 2024; Park et al.,, 2024). Elevated cytokines also disrupt endothelial homeostasis and increase vascular permeability (Park et al.,, 2024; Pope III et al.,, 2016; Briet et al.,, 2007). The sustained up-regulation of cytokines affects the hypothalamic thermoregulatory center, while impaired vascular permeability decreases the efficacy of peripheral vasodilatory response and increases the risk of heat-related hospital admissions (Vybíral et al.,, 2005).
@@ 4.4 Limitations and strengths
We acknowledge that our study has limitations. As with any case-crossover study, the results may not fully generalize to other populations or demographic groups; that is, the group of individuals who experienced a heat-related hospitalization is comprised of a small percentage of all Medicare FFS beneficiaries, and results drawn from this sample may not generalize to the full population. Furthermore, it is likely that PM2.5 partially mediates the relationship between temperature and various health conditions (Tran et al.,, 2023; Ham et al.,, 2024), including heat-related illnesses, so our estimate of the independent effect of temperature may not fully capture more complex pathways involving PM2.5. Consequently, interpreting temperature and PM2.5 as entirely independent exposures may be overly simple, as these exposures may influence each other’s effects on health outcomes in complex ways. We attempted to capture this with an interaction, but this may not sufficiently capture the true relationship between the exposures.
However, this study also has many strengths. We leverage a large and diverse sample of Medicare FFS beneficiaries, enabling robust estimates of risk across a range of environmental conditions. By applying a case-crossover design, the study effectively controls for time-invariant and slow-varying confounders at the individual level, allowing for more precise inferences of acute exposure effects. Moreover, our use of a flexible statistical model that incorporates both additive and interactive effects of temperature and PM2.5 provides valuable insights into the complex, real-world dynamics between these environmental exposures and their impacts on human health. Together, these strengths support the reliability of our findings and their relevance for guiding targeted public health interventions to mitigate the risks of heat-related hospitalizations in vulnerable populations.
@@ 4.5 Future work
As this study is the first to provide epidemiological evidence for synergism between the effects of temperature and PM2.5 on heat-related illness, additional work is needed to understand how these exposures interact. It has been well-established that the dangers of heat vary by geographic region and also time of year (Anderson and Bell,, 2011), since individuals adapt behaviorally and biologically to heat. An analysis that accounts for climate type or region may uncover differences in these effects by location. For example, the threat of PM2.5 may be greater in areas where fewer people have access to air conditioning and are thus more likely to be exposed to ambient air pollution on hot days. PM2.5 is also a diverse category of pollutants that includes a wide array of chemical components from various sources, and these components vary widely by location. Evidence suggests that certain components are more harmful than others (Hao et al.,, 2023; Masselot et al.,, 2022; Feng et al.,, 2024). Lastly, since this study focuses on older adults, there is a need to assess potential synergism among younger vulnerable individuals, such as children or adults who work outdoors. This study is a first step in understanding the complex synergism between the effects of temperature and PM2.5 on heat-related illness, and additional research is essential to understand the risk of simultaneous exposure.
Acknowledgments
This work was funded by the National Institutes of Health grants R01AG060232 and 1K01ES032458 and an Amazon Web Services grant on AI/ML for Identifying Social Determinants of Health.
Data availability and reproducibility
To protect the privacy of Medicare beneficiaries, our data usage agreement with CMS prohibits us from sharing patient data publicly. However, interested parties may purchase these data from CMS through the Research Data Assistance Center (ResDAC). The temperature and PM2.5 daily exposure data we used in our analysis are publicly available at https://prism.oregonstate.edu/ and https://doi.org/10.7910/DVN/58C6HG, respectively. Finally, our R analysis scripts are available at https://github.com/laurenmock/temp_pm25_heat-hosp.
Declaration of interests
Soko Setoguchi received research funding from Pfizer Inc., Pfizer Japan, Bristol Myers Squibb, and Daiichi Sankyo and served as a consultant for Pfizer Japan and Merck & Co., Inc.
Supplementary Materials
Appendix A Methods
A.1 ICD codes
@@
A.2 Main analysis details
@@
A.3 Sensitivity analysis details
As a sensitivity analysis, we also fit a model using a tensor product smooth to more flexibly capture the interaction between temperature and PM2.5. Tensor product smooths combine separate marginal smooth terms for each covariate, allowing each to have its own basis functions and smoothness penalties. Specifically, we implemented a t2-style tensor product smooth from the mgcv R package (Wood et al.,, 2013), which is compatible with Bayesian modeling in brms. Note that the alternative te-style tensor product smooths, which apply penalties differently, are currently not supported in brms. For further details and guidance on implementing tensor product smooths (t2 and te) using mgcv, see the tutorial by Pedersen et al., (2019).
A.4 OR and RERI estimation details
@@
Appendix B Additional results
| Main analysis (splines) | Sensitivity analysis (tensor) | |||
| Posterior | Posterior | |||
| Estimate (95% CI) | probability | Estimate (95% CI) | probability | |
| OR for heat (OR10) | 1.045 (1.026, 1.063) | <0.001 | 1.054 (1.027, 1.097) | <0.001 |
| OR for PM2.5 (OR01) | 1.014 (0.993, 1.037) | 0.098 | 1.032 (0.977, 1.101) | 0.116 |
| OR for heat & PM2.5 (OR11) | 1.091 (1.061, 1.124) | <0.001 | 1.106 (1.040, 1.194) | 0.002 |
| RERI | 0.032 (0.007, 0.057) | 0.006 | 0.019 (-0.028, 0.068) | 0.180 |
B.1 Sensitivity analysis results

### v2

@@ 3.2 Sensitivity analyses
First, we extended the temperature window from one day to three days and observed attenuated estimated associations between temperature and HRH with wider credible intervals relative to the main analysis (Figure 3 and Figures S3 and S4 in the Supplementary Material). Second, we modified the trimming threshold from the 95th percentile of the three-day PM2.5 distribution (18.4 ) to the 99th percentile (23.7 ). The findings indicate a much stronger positive association between PM2.5 and HRH than that seen in the main analysis, with narrow credible intervals that do not include one (Figure 3 and Figure S3 in the Supplementary Material). Finally, we developed an additional model using a tensor product to allow for a more flexible, nonlinear interaction between these exposures. This approach resulted in more linear estimates of the independent effects of temperature and PM2.5 compared with the other analyses (Figure S3 in the Supplementary Material). The estimated RERIs from all three analyses were positive, but the credible intervals from the sensitivity analyses with a longer temperature exposure and a tensor product included the null value of zero (Figure 3).
@@ 4.1 Summary and key findings
To our knowledge, this is the first study to investigate the combined (and synergistic) effects of acute exposure to high temperatures and elevated PM2.5 on HRHs. Comparing median and 95th percentile exposure levels, we found a positive association between temperature and HRH. The association between PM2.5 and HRH was also positive, although the credible interval covered the odds ratio scale null value of one. Finally, we identified evidence of an additive interaction between temperature and PM2.5.
Three sensitivity analyses demonstrate robustness in our findings and provide additional insights. Most notably, the analysis with a longer temperature exposure window indicated a weaker relationship between temperature and HRH, suggesting that the effects of temperature on HRH are more immediate. Additionally, the wider credible intervals from the tensor model reflect increased uncertainty in the estimates when allowing for greater flexibility in the interaction structure.
Our findings offer the first empirical, epidemiological evidence that PM2.5 may increase the risk of HRH. In contrast to prior work, which has largely framed temperature as a modifier of the health effects of air pollution, 13; 40; 41 our focus on explicit HRHs reveals that PM2.5 may also modify the health effects of heat.
@@ 4.2 Impact
Heat is the deadliest form of extreme weather in the United States, 42 and rates of heat-related illness continue to climb 9; 43 even though these illnesses are often preventable. 44; 45 Understanding how to protect vulnerable individuals, including older adults, from extreme heat events is a crucial component of climate change adaptation and an urgent public health priority. 9
@@
The National Weather Service issues localized heat alerts to initiate individual- and community-level preparations on unusually warm days. 46 Our findings indicate that additional caution is necessary on hot days when PM2.5 levels are high, and it may be useful to consider PM2.5 when issuing heat alerts. However, the efficacy of heat alerts alone in minimizing illness and death is uncertain. 47; 48 Previous research highlights that high-risk individuals must recognize their own vulnerabilities to heat, possess the knowledge and resources necessary to protect themselves, and take action to stay safe. 48 Interestingly, previous studies report that many older adults recognize age as a risk factor for heat-related illness but do not recognize high temperatures as a threat to their personal health. 49 Clinicians could play an important role in helping these individuals understand their risk. 44
@@
Our results also emphasize the importance of minimizing personal exposure to particulate matter, particularly on unusually warm days. For example, open windows may allow more ambient particulate matter into homes 50 and further increase the risk of heat-related illness. These findings also contribute to the growing body of evidence that PM2.5 is linked to a range of negative health effects, underscoring the importance of reducing ambient PM2.5, particularly with a warming planet.
4.3 Epidemiological evidence and biological plausibility of synergism
@@
This study supports mounting epidemiological evidence of synergism between temperature and PM2.5. Several studies have assessed effects on mortality; for example, a case-crossover study from Italy 12 identified positive but generally nonstatistically significant interactions between temperature and PM10, which includes both PM2.5 and larger particles, on mortality, and a meta-analysis by Li and colleagues 13 also reported that extreme temperatures modify the effect of PM10 on nonaccidental and cardiovascular mortality. An examination of PM2.5 and mortality in US cities 14 revealed a stronger association between long-term PM2.5 exposure and mortality in warmer cities. A smaller number have assessed hospitalization outcomes, including a recent study in New England that identified synergism between short-term exposures to temperature and PM2.5 for both respiratory and cardiac hospital admissions. 15 Our study contributes to this growing body of work and extends scientific understanding of synergism between these short-term exposures to HRHs, providing additional insight into the pathways between simultaneous temperature and air pollution exposures and morbidity and mortality.
@@
Furthermore, it is biologically plausible that PM2.5 may interfere with the body’s ability to respond to high temperatures. Simultaneous exposure to heat and PM2.5 can exacerbate heat-related health outcomes through shared pathophysiological mechanisms, including autonomic dysfunction and heightened inflammatory responses. Both heat and PM2.5 cause autonomic dysfunction by increasing sympathetic drive and suppressing parasympathetic activity. 51; 52 This imbalance leads to a 4%–7% reduction in heart rate variability and an increase in heart rate by 5–9 beats per minute. 53; 54; 55 In addition, blood pressure increases due to autonomic vasoconstriction from PM2.5 exposure, with each 10 increment correlating with a 9-mm Hg rise in systolic pressure. 56 Furthermore, increased sweating via sympathetic activation in response to heat elevates blood pressure by increasing plasma viscosity by 2% per C. 57 Thus, the combined cardiovascular stress disrupts hemodynamic regulation, compromising the body’s thermoregulatory function and exacerbating heat stress.
@@
Heat and PM2.5 can also increase systemic inflammatory responses by forming reactive oxygen species within the cells. The resulting oxidative stress activates redox-sensitive transcription factors, driving the production and release of pro-inflammatory mediators such as tumor necrosis factor-alpha, interleukin-6, and interleukin-1. 1; 58 A 10 increase in PM2.5 exposure corresponds to a five-fold rise in these cytokines, and a 2C elevation in ambient temperature can boost their expression by up to four-fold. 59; 60 Elevated cytokines also disrupt endothelial homeostasis and increase vascular permeability. 60; 61; 62 The sustained up-regulation of cytokines affects the hypothalamic thermoregulatory center, while impaired vascular permeability decreases the efficacy of peripheral vasodilatory response and increases the risk of heat-related hospital admissions. 63
@@ 4.4 Limitations and strengths
We acknowledge several limitations to our study. As with any case-crossover study, the results may not fully generalize to other populations or demographic groups; that is, the group of individuals who experienced an HRH is comprised of a small percentage of all Medicare FFS beneficiaries, and results drawn from this sample may not generalize to the overall US population. Furthermore, it is likely that PM2.5 partially mediates the relationship between temperature and various health conditions, 64; 65 including heat-related illness, so our estimate of the independent effect of temperature may not fully capture more complex pathways involving PM2.5. Consequently, interpreting temperature and PM2.5 as entirely independent exposures may be overly simple, as these exposures may influence each other’s effects on health outcomes in complex ways. We attempted to capture this with an interaction term, but this may not sufficiently capture the true relationship between the exposures. Moreover, due to the complexities of quantifying synergism between two continuous, time-varying exposures, our model did not capture daily exposures over time or interactions between lagged exposures that could be incorporated with a distributed lag model.
It is also possible that the observed relationship between temperature and PM2.5 is confounded by other pollutants such as ground-level ozone, which may be correlated with PM2.5 particularly in warm weather, 66 although previous studies have produced mixed results. 67 Finally, maximum daily temperature does not account for other meteorological variables such as humidity and wind speed, which may influence temperature perception; however, previous research indicates the choice of heat metric does not substantially alter epidemiological results. 68
This study also has several important strengths. We leverage a large and diverse sample of Medicare FFS beneficiaries, enabling robust estimates of risk across a range of environmental conditions. By applying a case-crossover design, the study effectively controls for confounders that are time-invariant or change slowly over time, allowing for more precise inferences of acute exposure effects. Moreover, our use of a flexible statistical model that incorporates both additive and interactive effects of temperature and PM2.5 provides valuable insights into the complex, real-world dynamics between these environmental exposures and their impacts on human health. Together, these strengths support the reliability of our findings and their relevance for guiding targeted public health interventions to mitigate the risks of HRH in vulnerable populations.
@@ 4.5 Future work
As this study is the first to provide epidemiological evidence for synergism between the effects of temperature and PM2.5 on heat-related illness, additional work is needed to understand how these exposures interact. It has been well-established that the dangers of heat vary by geographic region and also time of year, 69 since individuals adapt behaviorally and biologically to heat. An analysis that accounts for climate type or region may uncover differences in these effects by location. For example, the threat of PM2.5 may be greater in areas where fewer people have access to air conditioning and are thus more likely to be exposed to ambient air pollution on hot days. PM2.5 is also a diverse category of pollutants that includes a wide array of chemical components from various sources, and these components vary widely by location. Evidence suggests that certain components are more harmful than others. 70; 71; 72 Finally, since this study focuses on older adults, there is a need to assess potential synergism in younger vulnerable individuals, such as children or adults who work outdoors. This study is a first step in understanding the complex synergism between the effects of temperature and PM2.5 on heat-related illness, and additional research is essential to understand the risk of simultaneous exposure.
Acknowledgments: All computations presented here were supported by the Office of Research Computing at the Rutgers University Institute for Health, Health Care Policy and Aging.
Conflicts of interest: S.S. received research funding from Pfizer Inc., Pfizer Japan, Bristol Myers Squibb, and Daiichi Sankyo and served as a consultant for Pfizer Japan and Merck & Co., Inc. Other authors have no potential conflicts of interest to report.
Data availability: To protect the privacy of Medicare beneficiaries, our data usage agreement with the Centers for Medicare & Medicaid Services (CMS) prohibits us from sharing patient data publicly. However, interested parties may purchase these data from CMS through the Research Data Assistance Center (ResDAC). The temperature and PM2.5 daily exposure data we used in our analysis are publicly available at https://prism.oregonstate.edu/ and https://doi.org/10.7910/DVN/58C6HG, respectively. Finally, our R analysis scripts are available at https://github.com/laurenmock/temp_pm25_heat-hosp.
Source of funding: This work was funded by the National Institutes of Health grants R01AG060232 (principal investigator: S.S.) and 1K01ES032458 (principal investigator: R.N.) and an Amazon Web Services grant on AI/ML for Identifying Social Determinants of Health.
Supplementary materials for
“Excess risk of heat-related hospitalization associated with temperature and PM2.5 among older adults”
1 Methods
1.1 ICD codes
@@
1.2 Main analysis: natural cubic splines
@@
1.3 Sensitivity analysis: tensor product
As a sensitivity analysis, we also fit a model using a tensor product smooth to more flexibly capture the interaction between temperature and PM2.5. Tensor product smooths combine separate marginal smooth terms for each covariate, allowing each to have its own basis functions and smoothness penalties. Specifically, we implemented a t2-style tensor product smooth from the mgcv R package, 73 which is compatible with Bayesian modeling in brms. Note that the alternative te-style tensor product smooths, which apply penalties differently, are currently not supported in brms. For further details and guidance on implementing tensor product smooths (t2 and te) using mgcv, see the tutorial by Pedersen et al. 74
1.4 OR and RERI estimation details
@@
2 Exposure distributions
2.1 Sensitivity analysis results
We conducted three sensitivity analyses. First, we considered three-day average temperature exposures instead of same-day exposures. Second, we trimmed PM2.5 exposures at the 99th percentile instead of the 95th percentile. Lastly, we used a tensor product as opposed to natural cubic splines with an interaction term to capture non-linear interaction between exposures.
| Estimate (95% credible interval) | ||||
| Three-day | PM2.5 trimmed at | |||
| Main analysis | temperature | 99th percentile | Tensor product | |
| OR10 | 1.05 (1.03, 1.06) | 1.03 (1.01, 1.05) | 1.04 (1.02, 1.06) | 1.05 (1.03, 1.10) |
| OR01 | 1.01 (0.99, 1.04) | 1.09 (1.07, 1.12) | 1.04 (1.02, 1.06) | 1.03 (0.98, 1.10) |
| OR11 | 1.09 (1.06, 1.12) | 1.14 (1.11, 1.18) | 1.12 (1.09, 1.14) | 1.10 (1.04, 1.19) |
| RERI | 0.03 (0.01, 0.06) | 0.02 (-0.00, 0.05) | 0.03 (0.01, 0.05) | 0.02 (-0.03, 0.07) |